Several archeology students are unprepared for a surprise true/false test with 13 questions, and all of their answers are guesses. Find the mean for the number of correct answers for such students. (Round answer to one decimal place.) μ = Find the standard deviation for the number of correct answers for such students. (Round answer to two decimal places.) σ = Give the range for the usual number of correct answers. (Enter answer as an interval using square-brackets only with whole numbers.) usual values = Would it be unusual for a student to pass by guessing and getting at least 8 correct answers? no, it would not be unusual yes, it would be unusul?
Answer #1

Answer:

We have,

n=13

Probability of sucess = 1

That means,

P+q=1

Where,

p=0.5

q=1-p = 0.5

Mean = E(x)=np

=13*0.5

Mean = 6.5

Standard deviation = (npq)^(1/2)

= (13*0.5*.0.5)^(1/2)

Standard deviation = 1.802